[Ag(NH3)2]+ (aq) + e- \rightarrow Ag(s) + 2 NH3 (aq) E^{\circ}= +0.373V

Use the standard redox potential of silver below to calculate the stability constant of the silver diamine complex.

Ag+ + e- \rightarrow Ag(s) E^{\circ}?= +0.7996V

Would you expect the redox potential to increase or decrease if NH3 is replaced with pyridine? Explain your answer
Answer #1

Ag + + e-\rightarrow Ag (S) E^{\circ}​= +0.7996V

[Ag(NH3)2+] (aq) + e- \rightarrow Ag(s) + 2 NH3 (aq)    E^{\circ}= +0.373V

Hence overall reaction is

Ag + +  2 NH3 (aq) \rightarrow [Ag(NH3)2+] (aq)

Eo Cell = Eo cathode - Eo anode = 0.7996V- 0.373V= 0.4266V

Further Eo = 0.0591 log k

log k = 0.4266V/ 0.0591 = 7.22 , therefore

Staility constant k= 1.6 x 107